OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10

OpenStudy (anonymous):

\[ 2(x − 4)=20\iff x-4=10\iff x=14\]

Parth (parthkohli):

You should equate the values. \( \color{Black}{\Rightarrow 20 = 2(x - 4) }\) First step would be to distribute. \( \color{Black}{\Rightarrow 20 = 2x - 8}\) Please tell me that you know how to solve equations :P